Determine whether the given linear equations are parallel, perpendicular, or neither.
y = - 5x - 8
y = 5x +11
Please show steps. Thank you.

Answer by pragya.ban@gmail.com(32)

Determine whether the given linear equations are parallel, perpendicular, or neither.
1) y = - 5x - 8
2) y = 5x +11

First of equations are parallel to one another if they share the same slope. (ex: y= 2x+3 and y= 2x+4) Equations are perpendicular if they are a negative reciprocal of each other. (ex: y= 3x+9 and y= -1/3x+5) A way to see if two equations are perpendicular is to multiply the slopes and if they equal to -1 then they are perpendicular.

For equation number 1, the slope is -5x, and for equation number 2 the slope is 5x. Based on what we just described above, the two slopes are not the same (-5 is not equal to 5) and they are not negative reciprocals of each other either (-5*5 does not equal to -1).

Therefore these two linear equations are neither parallel nor perpendicular.